# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ 3.5 Impact of gaps in the galaxy catalog
A potentially problematic case that can arise in observations of the GaLCC is when the galaxy detection method employed results in gaps in the galaxies identified around the line of sight. This can arise e.g.Ā if a set of different emission lines is employed to identify galaxies at different redshifts (i.e.Ā at different positions along the line of sight) because of the finite spectral coverage of the instrument employed. We test the impact of such artefact in the galaxy selection by manually removing galaxies in specific (small) redshift windows along a sightline. Specifically, we sample the length of each such gap from a uniform distribution in the range , where is the length of each spectrum and is the number of gaps in each spectrum. This ensures that, on average, of the galaxies are masked out. We produce 50 random realizations of gaps and show in Fig.Ā 11 the envelope of all resulting GaLCC for varying number of gaps. We find that fewer (and hence longer) gaps more strongly impact the inferred GaLCC than more shorter ones. We also test the impact of fixing the gap length and changing the gap number, that however shows results fully consistent with the changing number of galaxies selected (since in this case the average masked-out region increases with gap number).
@@ 3.6 Lightcone effects
In order to test the impact of lightcone effects on the GaLCC, we have included them in our synthetic Ly forest spectra (using a step-wise constant approximation). This allows us to explore the impact of the redshift window employed in observation of this quantity. We show in Fig.Ā 12 the resulting cross-correlation for a redshift window centred at and of varying width (solid colored lines). For reference, we also show the cross-correlation computed using our coeval (i.e. without any lightcone effect) at (dashed, dotted and dot-dashed black lines, respectively). Note that, since for these coeval sightlines the entire synthetic spectrum lies at a fixed redshift, the total path length covered by these spectra is larger (for the same number of sightlines considered, as it is the case here). Therefore, we expect small differences in the cross-correlation computed from the lightcone spectra fixing and the one computed from the coeval spectra. However, if we were to match the spectrum path (in both length and position in the simulation box), the result from the lightcone spectra would be by construction identical to the one from the coeval sightlines.
@@
We have chosen to focus on for a number of reasons. First, observations of the GaLCC (e.g.Ā Kakiichi etĀ al., 2018; Meyer etĀ al., 2019, 2020; Kashino etĀ al., 2023) are at or close to this redshift. Second, it approximately marks the beginning of the tail end of reionization, characterised by a rapid evolution of the volume-averaged HāI fraction as the ionised regions overlap. This makes such period especially interesting for a number of reasons: (i) it induces a rapid evolution of the IGM global quantities, enabling an accurate timing of the EoR and, in turn, tight constraints on the source properties; (ii) the flux in the Ly forest is non-negligible (unlike at e.g.Ā ) but the fluctuations in the UV field are still strong enough to leave imprints in the galaxy-Ly correlation (unlike at e.g.Ā Meyer etĀ al., 2020). Third, since our simulations end at , choosing allow us to study the impact of redshift windows up to without running into the issue that no more outputs are available.
@@
Fig.Ā 12 shows that the impact of lightcone effects is negligible for , essentially only affecting the smoothness of the curve thanks to improved statistics when allowing for larger redshift bins, that therefore include more galaxies. However, for larger the lightcone effects become relevant. In particular, the amplitude of the excess is reduced and the position of the peak moves to larger distances. This occurs because the signal becomes dominated by the lower-redshift part of the redshift window as a consequence of the larger number of galaxies (and therefore pixel-galaxy pairs) found in it. In fact, the curve approaches more and more the dot-dashed lines showing the cross-correlation signal recorded at the lowest redshift encapsulated in the largest redshift window.
@@
This has important implications for studies that cover broad redshift windows, es e.g.Ā Meyer etĀ al. (2019, 2020). In particular, it implies that in order to properly compare such observations with theoretical and numerical predictions, lightcone-like sightlines are needed. This requirement was not met by previous studies (e.g.Ā Garaldi etĀ al., 2022), but our result demonstrate how this is necessary for an accurate comparison. Alternatively, an approximate workaround is to compare observation to prediction made at an effective redshift encapsulating the biased mixing of signals coming from different redshift ranges. In order to investigate this possibility, we show in Fig.Ā 13 the cross-correlation signal computed from the lightcone los using (dashed blue curve, same as in Fig.Ā 12) alongside the correlation signal computed at fixed time (solid lines) for a variety of redshift (indicated by the colorbar). It can be seen that the cross-correlation computed from coeval sightlines at almost perfectly coincides with the one on obtained from the lightcone los using the redshift window . From the previous discussion, it follows that a good candidate for this is the median redshift of the galaxies used to compute the cross correlation. We have tested that this is indeed the case even for the largest values of probed (, see also AppendixĀ C for more details).
@@
Finally, we note that in Garaldi etĀ al. (2022) the predictions of the thesan simulations were compared the data collected by Meyer etĀ al. (2019), that compound observations over the broad redshift window . In that study, the simulations output at were used, as this is very close to the central redshift of the observed data. In light of the results discussed here, it would have been more appropriate to compare to a lower redshift, that however is unfortunately not available for the thesan simulations. Nevertheless, extrapolating the redshift evolution found in thesan (Fig.Ā 16 of Garaldi etĀ al., 2022), we can speculate that using a lower-redshift output for the comparison would worsen the tension found between the simulated and observed cross-correlation.
@@ 4.1 Connection to the local overdensity
As discussed in Sec.Ā 3, the selection of galaxies to observe has an impact on the resulting cross-correlation signal. Here we aim to address the questions: What do we learn by looking at different galaxy populations? Is there an optimal galaxy sample? In Garaldi etĀ al. (2022) it was already shown that the amplitude of the flux enhancement depends on the galaxy used to compute the cross-correlation. In particular, more biased tracers (i.e.Ā more massive galaxies and haloes) yield a stronger signal. At the same time, it was also shown that, once the reionization history is accounted for, models where small galaxies dominate the photon budget produce approximately the same GaLCC as models dominated by large galaxies. To elucidate such counter-intuitive result, here we supplement that study with by investigating the impact of the local overdensity of the galaxies selected. In fact, a number of relevant physical processes relevant for the cross-correlation signal investigated here are sensitive to the local overdensity. First, higher local densities imply larger recombination rates that can suppress the flux in the Ly forest. Second, overdense regions hosts more galaxies, whose radiation output combines to create larger ionised bubbles and stronger radiation fields. Finally, larger overdensities are more likely to host massive galaxies.
@@
For the purpose of this study, we define the local overdensity of a galaxy () as in Neyer etĀ al. (2023). In practice, we first smooth the density field with a Gaussian filter with standard deviation 1 cMpc. We then define , where is the average gas density in the Universe at the given redshift and is the smoothed density field. Equipped with this definition, we compute the cross-correlation signal in bins of stellar mass and local overdensity. More specifically, the former are delimited by and the latter by . We show the resulting correlation function in Fig.Ā 14, where the different panels correspond to different bins in stellar mass (reported in the bottom right corner of each panel). The lines within each panel show the result of isolating different local overdensities (shades of green) and of considering all galaxies together (black line). The effect of the local overdensity can be quite strong. Larger increases the amplitude of the measured GaLCC, but does not affect its shape. This has two important implications, that we describe next.
@@
The finding described above has important implications for the interpretation of the observed signal. In fact, a random sightline to a QSO will cross (or pass in the proximity of) regions of different overdensity, each one imprinting its own intrinsically different signature, contributing to the āstructure formation noiseā described in Sec.Ā 3.1. This is not problematic if the cross-correlation is computed over a sufficiently-large number of sightlines, but can bias the results if this number is limited. In this case, the impact of a sightlines sampling an extreme overdensity can be very large, which renders it difficult to connect the observed signal to the physical properties of the Universe. Similarly, if the galaxy detection method employed results in galaxies spatially clustered around few specific positions of the spectrum (e.g.Ā when observations are sensitive to just extremely bright objects, which tend to be clumped in large overdensities), the final result will be driven by one or few overdensitied where such objects reside.
@@ 4.1.1 Insensitivity to source bias
The results described above are linked to an intriguing finding of Garaldi etĀ al. (2022), namely that once the different reionization histories are factored out, the cross-correlation signal does not depend on the sources of reionization. In fact, in their Fig.Ā 17, the signal from the thesan-low-2 and thesan-high-2 simulations (which emit ionizing photons only from galaxies more and less massive than , respectively) are essentially indistinguishable from the one stemming from their standard runs. This appears at odds with their result that galaxies produce a modulation of the Ly flux in their local environment that scales with their mass. These two apparently-contradicting statements can be reconciled once the role of overdensity is taken into account. For example, in the case of thesan-low-2, although only small galaxies are emitting ionizing photons, they are still preferentially found in overdensities, where also larger object tend to reside. This, combined with their larger photon output (i.e.Ā higher escape fraction, needed to achieve a somewhat similar reionization history) entails that overall the contribution to the GaLCC still remains dominated by large conglomerates of small galaxies, that essentially act as rare bright sources in this model.
@@
In order to show this explicitly, in Fig.Ā 15 we display the cross-correlation signal computed at for all galaxies (solid blue line), for small galaxies (i.e.Ā , dashed red line), for small galaxies with local overdensity (dotted yellow line), for large galaxies (i.e.Ā , solid green line) and for large galaxies with local overdensity (dot-dashed purple line). A first striking feature is that the exclusion of big galaxies does not affect the correlation signal at all (compare the solid blue and dashed red lines), clarifying that the impact of large objects in term of ionizing photons is negligible (as expected from the ionizing photons budget discussed in Kannan etĀ al., 2022a). However, selecting large galaxies to compute the cross-correlation (as done in Garaldi etĀ al., 2022) means selecting preferentially overdense regions (because of the larger bias associated to such objects). The impact of such selection can be seen comparing the dashed red and dotted yellow line. Both curves are computed only for small galaxies, but the latter additionally selects only objects with , resulting in a significantly boosted signal. In fact, such boost is very similar to the one obtained by selecting large galaxies only (solid green line). Symmetrically, removing galaxies in large overdensities from the sample of massive objects (dot-dashed purple line, i.e.Ā artificiallty removing the preference of large galaxies to reside in overdense regions) brings the signal down to a level very similar to the one obtained including all galaxies.
@@
From Fig.Ā 15 and the related discussion above, it appears that the main driver of the flux enhancement at intermediate scales are overdense regions and not bright sources. Naturally, these two are related in the CDM cosmology, since larger/brighter sources are more biased than smaller/fainter ones and preferentially reside in overdense regions. However, our results show that it is the clustering of multiple faint sources (which occurs primarily in overdense regions) to produce the observed signal, not individual/rare bright sources, as a consequence of the complexity of the Ly signal and of the additional recombinations occurring in overdense regions counteracting the additional ionizing photons emitted by the rare bright sources (which are sub-dominant with respect to the one emitted by smaller galaxies, see e.g.Ā Rosdahl etĀ al., 2022; Kostyuk etĀ al., 2023; Yeh etĀ al., 2023). This finding reconciles the apparently contradicting results that the ionizing photons budget is dominated by small galaxies, that the GaLCC is enhanced around bright objects and that restricting the ionizing photons production to the largest/smallest galaxies does not affect the signal (once the different reionization histories are factored out).
@@ 4.2 Dependence on the sightline effective optical depth
The IGM ionised fraction along the line of sight is expected to affect the inferred GaLCC. In fact, more neutral sightlines are more opaque to Ly photons (i.e.Ā their effective optical depth is larger), decreasing the total transmitted flux. For this reason, following Garaldi etĀ al. (2019) the GaLCC is typically normalised by the average transmitted flux, in order to factor out sightline-to-sightline variations in the average density (and, hence, neutral hydrogen fraction). However, this can create artefacts that need to be understood. To clarify this, we present in Fig.Ā 16 the GaLCC computed for all sightlines in thesan-1 (black solid line), using only the 100 most opaque ones (purple dashed line) and using only the 100 most transparent ones (yellow dashed line). Comparing the dashed curves it appears clearly that the peak in transmissivity is larger in amplitude and located at smaller distances from galaxies when only the most opaque sightlines are considered. This might appear counter-intuitive given the interpretation of this as excess ionization (compared to the average) in the proximity of galaxies. However, this is a spurious effect of the normalization chosen. In practice, since each line of sight is not a fair sample of the distribution of densities and ionised fractions in the Universe, the normalization is larger (smaller) for the most transparent (opaque) sightlines, boosting the GaLCC computed from the latter. This is shown explicitly in the Figure by the shaded regions (to be read against the right-hand-side vertical axis) that report the average transmitted flux as function of distance (without normalization) for the same two sightlines samples. It now appears clear that transparent lines of sight show approximately 4 times more Ly flux than the opaque ones. At the same time, the peak is located to smaller distances from galaxies, likely as a consequence of the increased absorption of photons due to the more neutral environment, which constrain this ionizing radiation-driven proximity effect to smaller distances. We discuss this in more details next.

### v2

@@ 3.5 Impact of gaps in the galaxy catalog and spectrum
A potentially problematic case that can arise in observations of the GaLCC is when the galaxy detection method employed results in gaps in the galaxies identified around the line of sight. This can arise e.g. if a set of different emission lines is employed to identify galaxies at different redshifts (i.e. at different positions along the line of sight) because of the finite spectral coverage of the instrument employed. We test the impact of such an artefact in the galaxy selection by manually removing galaxies in specific (small) redshift windows along a sightline. Specifically, we sample the length of each such gap from a uniform distribution in the range , where is the length of each spectrum and is the number of gaps in each spectrum. This ensures that, on average, of the galaxies are masked out. We produce 50 random realizations of gaps and show in the top panel of Fig. 10 the envelope of all resulting GaLCC for varying number of gaps. We find that fewer (and hence longer) gaps more strongly impact the inferred GaLCC than more shorter ones. We also test the impact of fixing the gap length and changing the gap number. This shows results fully consistent with the changing number of galaxies selected (since in this case the average masked-out region increases with gap number).
We also test the case in which it is instead the spectrum to have gaps, due for example to removal of foreground contamination. Since in this case it is more likely to have a large number of small gaps rather than one or few very large ones, we instead quantify the impact of masking different fractions of the spectrum, sampling the gaps length from the range (note that, unlike in the galaxy gaps case, this does not depends on the gap number). We show in the bottom panel of Fig. 10 that the GaLCC is very resilient to this type of masking, likely because the number of pixels in the spectrum is sufficiently high so that, even when a large fraction of them is removed, the underlying signal can be recovered with small uncertainty. We have also checked that, as expected, performing the same gap number-dependent sampling of the gap lengths as we have done for the galaxy catalog results in almost identical results.
@@ 3.6 Lightcone effects
In order to test the impact of lightcone effects on the GaLCC, we have included them in our synthetic Ly forest spectra and galaxy catalog (using a step-wise constant approximation). This allows us to explore the impact of the redshift window employed in observation of this quantity. We show in Fig. 11 the resulting cross-correlation for a redshift window centred at and of varying width (solid colored lines). For reference, we also show the cross-correlation computed using our coeval (i.e. without any lightcone effect) at (dashed, dotted and dot-dashed black lines, respectively). Note that, since for these coeval sightlines the entire synthetic spectrum lies at a fixed redshift, the total path length covered by these spectra is larger (for the same number of sightlines considered, as it is the case here). Therefore, we expect small differences in the cross-correlation computed from the lightcone spectra fixing 222Since for the spectrum length is vanishing, we have approximated this by considering only the part of each spectrum associated to , which is non-vanishing because of the stepwise-constant approximation used in producing the lightcone-like spectra (see Sec. 2.2). In practice, thanks to our high time cadence of snapshots, this is equivalent to . and the one computed from the coeval spectra. However, if we were to match the spectrum path (in both length and position in the simulation box), the result from the lightcone spectra would be by construction identical to the one from the coeval sightlines.
@@
We have chosen to focus on for a number of reasons. First, observations of the GaLCC (e.g. Kakiichi et al., 2018; Meyer et al., 2019, 2020; Kashino et al., 2023) are at or close to this redshift. Second, it approximately marks the beginning of the tail end of reionization, characterised by a rapid evolution of the volume-averaged H I fraction as the ionized regions overlap. This makes such period especially interesting for a number of reasons: (i) it induces a rapid evolution of the IGM global quantities, enabling an accurate timing of the EoR and, in turn, tight constraints on the source properties; (ii) the flux in the Ly forest is non-negligible (unlike at e.g. ) but the fluctuations in the UV field are still strong enough to leave imprints in the galaxy–Ly correlation (unlike at e.g. Meyer et al., 2020). Third, since our simulations end at , choosing allow us to study the impact of redshift windows up to without running into the issue that no more outputs are available.
@@
Fig. 11 shows that the impact of lightcone effects is negligible for , essentially only affecting the smoothness of the curve thanks to improved statistics when allowing for larger redshift bins, that therefore include more galaxies. However, for larger the lightcone effects become relevant. In particular, the amplitude of the excess is reduced and the position of the peak moves to larger distances. This occurs because the signal becomes dominated by the lower-redshift part of the redshift window as a consequence of the larger number of galaxies (and therefore pixel-galaxy pairs) found in it. In fact, the curve approaches more and more the dot-dashed lines showing the cross-correlation signal recorded at the lowest redshift encapsulated in the largest redshift window.
@@
This has important implications for studies that cover broad redshift windows, e.g. Meyer et al. (2019, 2020). In particular, it implies that in order to properly compare such observations with theoretical and numerical predictions, lightcone-like sightlines are needed. This requirement was not met by previous studies (e.g. Garaldi et al., 2022), but our results demonstrate how this is necessary for an accurate comparison. Alternatively, an approximate workaround is to compare observation to prediction made at an effective redshift encapsulating the biased mixing of signals coming from different redshift ranges. In order to investigate this possibility, we show in Fig. 12 the cross-correlation signal computed from the lightcone LOS using (dashed blue curve, same as in Fig. 11) alongside the correlation signal computed at fixed time (solid lines) for a variety of redshift (indicated by the colorbar). It can be seen that the cross-correlation computed from coeval sightlines at almost perfectly coincides with the one on obtained from the lightcone LOS using the redshift window . From the previous discussion, it follows that a good candidate for this is the median redshift of the galaxies used to compute the cross correlation. We have tested that this is indeed the case even for the largest values of probed (, see also Appendix D for more details).
@@
Finally, we note that in Garaldi et al. (2022) the predictions of the thesan simulations were compared the data collected by Meyer et al. (2019), that compound observations over the broad redshift window . In that study, the simulations output at were used, as this is very close to the central redshift of the observed data. In light of the results discussed here, it would have been more appropriate to compare to a lower redshift, that however is unfortunately not available for the thesan simulations. Nevertheless, extrapolating the redshift evolution found in thesan (Fig. 16 of Garaldi et al., 2022), we can speculate that using a lower-redshift output for the comparison would worsen the tension found between the simulated and observed cross-correlation, requiring an even later end of reionization to match the observed signal.
@@ 4 Physical processes connected to the GaLCC
The leading interpretation of the GaLCC paints a picture where the observed signal arises from two opposite effects, each dominating at different scales around galaxies. The enhanced ionizing photon density provided by (large) galaxies boosts the transmission of Ly in their surroundings. This transverse proximity effect can be detected statistically by looking at the excess Ly transmission around galaxies. At the same time, the overdensity where galaxies reside suppress the Ly transmission, thanks to their increased density and, hence, boosted H II recombination rate. This is the dominant effect at the smallest scales, while the aforementioned radiation proximity effect takes over at intermediate ones. In Fig. 13 we visually show this by reporting in the top panel the GaLCC computed from thesan-1 at and using galaxies with stellar mass (black solid line). We mark the radii corresponding to , , , and Mpc with light grey vertical solid lines. In the middle and bottom panels, we show the stacked ionizing photon density and gas density around the same galaxies used in the computation of the GaLCC. In order to compute these quantities, we extract gas properties in a cubic region with side Mpc and centered on each galaxy in the sample. We then stack these regions and finally project the resulting cube along one axis for visualisation purposes. In the middle and bottom panel, we mark with dashed circles the same radii as above. It can clearly be seen that the central regions (approximately within - Mpc) are dominated by high densities, which suppress the Ly transmission thanks to the boosted recombination rate suppressing the ionized fraction. At intermediate scales () the density has already fallen to the background level, but the ionizing photons density is still largely above the one observed farther away from the center. The latter enhances the Ly transmission at such scales, resulting in a broad peak in the GaLCC.
@@ 4.1 Connection to the local overdensity
As discussed in Sec. 3, the selection of galaxies to observe has an impact on the resulting cross-correlation signal. Here we aim to address the questions: What do we learn by looking at different galaxy populations? Is there an optimal galaxy sample? In Garaldi et al. (2022) it was already shown that the amplitude of the flux enhancement depends on the galaxy used to compute the cross-correlation. In particular, more biased tracers (i.e. more massive galaxies and haloes) yield a stronger signal. At the same time, it was also shown that, once the reionization history is accounted for, models where small galaxies dominate the photon budget produce approximately the same GaLCC as models dominated by large galaxies. To elucidate such a counter-intuitive result, here we supplement that study with by investigating the impact of the local overdensity of the galaxies selected. In fact, a number of relevant physical processes relevant for the cross-correlation signal investigated here are sensitive to the local overdensity. First, higher local densities imply larger recombination rates that can suppress the flux in the Ly forest. Second, overdense regions hosts more galaxies, whose radiation output combines to create larger ionized bubbles and stronger radiation fields. Finally, larger overdensities are more likely to host massive galaxies.
@@
For the purpose of this study, we define the local overdensity of a galaxy () as in Neyer et al. (2024). In practice, we first smooth the density field with a Gaussian filter with standard deviation 1 cMpc. We then define , where is the average gas density in the Universe at the given redshift and is the smoothed density field. Equipped with this definition, we compute the cross-correlation signal in bins of stellar mass and local overdensity. More specifically, the former are delimited by and the latter by . We show the resulting correlation function in Fig. 14, where the different panels correspond to different bins in stellar mass (reported in the bottom right corner of each panel). The lines within each panel show the result of isolating different local overdensities (shades of green) and of considering all galaxies together (black line). The effect of the local overdensity can be quite strong. Larger increases the amplitude of the measured GaLCC, but does not affect its shape.
@@
The finding described above has important implications for the interpretation of the observed signal. In fact, a random sightline to a QSO will cross (or pass in the proximity of) regions of different overdensity, each one imprinting its own intrinsically different signature, contributing to the ‘structure formation noise’ described in Sec. 3.1. This is not problematic if the cross-correlation is computed over a sufficiently-large number of sightlines, but can bias the results if this number is limited. In this case, the impact of a sightlines sampling an extreme overdensity can be very large, which renders it difficult to connect the observed signal to the physical properties of the Universe. Similarly, if the galaxy detection method employed results in galaxies spatially clustered around few specific positions of the spectrum (e.g. when observations are sensitive to just extremely bright objects, which tend to be clumped in large overdensities), the final result will be driven by one or few overdensitied where such objects reside.
@@ 4.1.1 Insensitivity to source bias
The results described above are linked to an intriguing finding of Garaldi et al. (2022), namely that once the different reionization histories are factored out, the cross-correlation signal does not depend on the sources of reionization. In fact, in their Fig. 17, the signal from the thesan-low-2 and thesan-high-2 simulations (which emit ionizing photons only from galaxies more and less massive than , respectively) are essentially indistinguishable from the one stemming from their standard runs. This appears at odds with their result that galaxies produce a modulation of the Ly transmission in their local environment that scales with their mass. These two apparently-contradicting statements can be reconciled once the role of overdensity is taken into account. For example, in the case of thesan-low-2, although only small galaxies are emitting ionizing photons, they are still preferentially found in overdensities, where also larger objects tend to reside. This, combined with their larger photon output (i.e. higher escape fraction, needed to achieve a somewhat similar reionization history) entails that overall the contribution to the GaLCC still remains dominated by large conglomerates of small galaxies, that essentially act as rare bright sources in this model.
@@
In order to show this explicitly, in Fig. 15 we display the cross-correlation signal computed at for all galaxies (solid blue line), for small galaxies (i.e. , dashed red line), for small galaxies with local overdensity (dotted yellow line), for large galaxies (i.e. , solid green line) and for large galaxies with local overdensity (dot-dashed purple line). A first striking feature is that the exclusion of big galaxies does not affect the correlation signal at all (compare the solid blue and dashed red lines), clarifying that the impact of large objects in term of ionizing photons is negligible (as expected from the ionizing photons budget discussed in Kannan et al., 2022a). However, selecting large galaxies to compute the cross-correlation (as done in Garaldi et al., 2022) means selecting preferentially overdense regions (because of the larger bias associated to such objects). The impact of such selection can be seen comparing the dashed red and dotted yellow line. Both curves are computed only for small galaxies, but the latter additionally selects only objects with , resulting in a significantly boosted signal. In fact, such boost is very similar to the one obtained by selecting large galaxies only (solid green line). Symmetrically, removing galaxies in large overdensities from the sample of massive objects (dot-dashed purple line, i.e. artificiallty removing the preference of large galaxies to reside in overdense regions) brings the signal down to a level very similar to the one obtained including all galaxies.
@@
From Fig. 15 and the related discussion above, it appears that the main driver of the transmissivity enhancement at intermediate scales are overdense regions and not bright sources. Naturally, these two are related in the CDM cosmology, since larger/brighter sources are more biased than smaller/fainter ones and preferentially reside in overdense regions. However, our results show that it is the clustering of multiple faint sources (which occurs primarily in overdense regions) to produce the observed signal, not individual/rare bright sources, as a consequence of the complexity of the Ly signal and of the additional recombinations occurring in overdense regions counteracting the additional ionizing photons emitted by the rare bright sources (which are sub-dominant with respect to the one emitted by smaller galaxies, see e.g. Rosdahl et al., 2022; Kostyuk et al., 2023; Yeh et al., 2023). This finding reconciles the apparently contradicting results that the ionizing photons budget is dominated by small galaxies, that the GaLCC is enhanced around bright objects and that restricting the ionizing photons production to the largest/smallest galaxies does not affect the signal (once the different reionization histories are factored out).
@@ 4.2 Dependence on the sightline effective optical depth
The IGM ionized fraction along the line of sight is expected to affect the inferred GaLCC. In fact, more neutral sightlines are more opaque to Ly photons (i.e. their effective optical depth is larger), decreasing the total transmitted flux. For this reason, following Garaldi et al. (2019) the GaLCC is typically normalised by the average transmitted flux, in order to factor out sightline-to-sightline variations in the average density (and, hence, neutral hydrogen fraction). However, the choice of whether to compute the average transmitted flux from the observed spectra sample or to use some independent measurement can create artifacts that need to be understood. We stress that, as outlined in Sec. 2.3, this distinction is practically non-existent for our work, by virtue of how the number and production method of our synthetic spectra.
We present in Fig. 16 the GaLCC computed for all sightlines in thesan-1 (black solid line), using only the 100 most opaque ones (purple dashed line) and using only the 100 most transparent ones (yellow dashed line) and normalised by the average flux in the sample. Comparing the dashed curves it appears clearly that the peak in transmissivity is larger in amplitude and located at smaller distances from galaxies when only the most opaque sightlines are considered. This might appear counter-intuitive given the interpretation of this as excess ionization (compared to the average) in the proximity of galaxies. However, this is a spurious effect of the normalization chosen. In practice, since each line of sight is not a fair sample of the distribution of densities and ionized fractions in the Universe, the normalization is larger (smaller) for the most transparent (opaque) sightlines, boosting the GaLCC computed from the latter. This is shown explicitly in the Figure by the shaded regions that report the average transmitted flux as function of distance for the same two sightlines samples but now normalized to the average transmitted flux in the entire simulation (and therefore sharing the same normalization). It now appears clear that transparent lines of sight show approximately 4 times more Ly transmission than the opaque ones. At the same time, the peak is located to smaller distances from galaxies, likely as a consequence of the increased absorption of photons due to the more neutral environment, which constrain this ionizing radiation-driven proximity effect to smaller distances. We discuss this in more detail next.